OpenStudy (sidsiddhartha):

an ellipse of eccentricity [2sqrt(2)]/3 is inscribed in a circle and a point within the circle is chosen at random.what is the probability that the point lies outside the ellipse

OpenStudy (amistre64):

do you recall what the eccentricity measures?

OpenStudy (amistre64):

e = c/a .... focal length in terms of vertex length

OpenStudy (amistre64):



OpenStudy (sidsiddhartha):

yeah its\[\sqrt{(1-\frac{ b ^{2} }{ a ^{2}}}\]

OpenStudy (amistre64):

the area of the circle if pi a^2

OpenStudy (sidsiddhartha):

yeah if a=b then its pia^2 then?

OpenStudy (amistre64):

if a=b its a circle, and e=1

OpenStudy (amistre64):

the circle that the ellipse is drawn inside of has a radius of a

OpenStudy (sidsiddhartha):

yeah ok

OpenStudy (amistre64):



OpenStudy (amistre64):

P(inside ellipse) = area ellipse/ area circle so 1- that is outside the ellipse